# On the solvable-unsolvable transition due to noise-induced chaos in digital memcomputing

### v1

@@ Abstract
keywords:
Chaos, nonlinear dynamical systems, computing technology, digital memcomputing
@@
[1]organization=Department of Physics and Astronomy, University of South Carolina, city=Columbia, SC, postcode=29208, country=USA
[…]
[2]organization=École Polytechnique Fédérale de Lausanne, city=Lausanne, postcode=1015, country=Switzerland
[…]
[3]organization=Department of Physics, University of California, San Diego, city=La Jolla, CA, postcode=92093-0319, country=USA
1 Introduction
In the field of computing, a new class of non-linear (classical) dynamical systems has been recently introduced, designed to solve complex combinatorial optimization problems [1]. These dynamical systems have been named digital memcomputing machines (DMMs) in view of the fact that they map a finite string of symbols to a finite string of symbols while coupling two types of degrees of freedom: “fast ones” representing the variables of the problem to solve, and “slow ones” representing memory degrees of freedom [2]. Since DMMs are classical systems, their ordinary differential equations (ODEs) can be efficiently simulated on our modern computers. In fact, most of the studies concerning DMMs so far have been performed by simulating their ODEs [3, 4, 5, 6, 7, 8, 9] [^1] Also see, e.g., case studies in [2] or those reported by the company MemComputing, Inc.: www.memcpu.com., with a few implementations on small and large FPGA boards [10, 11].
Despite much work in this fledgling field has focused on understanding the topological aspects of these dynamical systems [12, 13], less attention has been devoted to the role of noise, whether in the form of numerical errors introduced in the discretization of the corresponding ODEs, or the physical noise that would naturally appear in a hardware realization of these machines. Numerical noise can be increased during the time discretization process, by increasing the integration time step, resulting in discrete approximations that alter the path of the system in phase space. Physical noise occurs due to variations in the environment or control parameters, adding complexity and unpredictability. Although there have been multiple studies investigating noise effects in discrete maps and Hamiltonian systems [14, 15], the influence of noise on the transition to chaos within continuous-time dynamical systems remains less examined [16]. Earlier research was predominantly focused on either systems with a high dimensionality [17, 18] or pronounced non-linearities [19, 20, 21].
Here, we examine the noise-induced transition from a solvable to an unsolvable phase in prototypical DMMs as those introduced in [7]. Such DMMs have been designed to solve 3-SAT instances (a collection of logical disjunctions of exactly three variables or their negations, related to each other by logical conjunctions), a well-known combinatorial problem [22]. These DMMs solve these instances efficiently thanks to the emergence of dynamical long-range order (DLRO) [13] and the fact that, if properly designed, they admit only saddle points and equilibria corresponding to the solution(s) of the problem instance. No other local minima are present in the phase space.
In the language of field theory, the DLRO of DMMs is expressed by the collective behavior of specific trajectories in the phase space, called instantons, or avalanches [13], which connect critical points of the flow vector field of the dynamics with decreasing index, namely from saddle points with a certain number of unstable directions to saddle points with less unstable directions. We refer to the book [2] for a thorough explanation of these statements and to Ref. [7] for their mathematical demonstration. The very fact that DMMs employ critical points (saddle points) and trajectories between them (instantons) make them topologically robust, in the sense that their dynamics are fairly insensitive to perturbations. However, we expect this robustness to reach a limit as a function of the noise strength introduced during dynamics, whether numerical or physical.
The structure of this paper is as follows. Sec. 2 describes the methods employed in this investigation. In Sec. 3, we detail the modeling results, beginning with the case of numerical noise (Sec. 3.1) and subsequently addressing the case of physical noise (Sec. 3.2). We discuss our results in Sec. 4, and conclude in Sec. 5.
2 Methods
2.1 DMMs equations for 3-SAT
Let us consider a Boolean satisfiability problem referred to as 3-SAT, which comprises $N$ Boolean variables and $M$ clauses. Each clause is a disjunction of three literals (variables or their negation). A solution to the 3-SAT instance involves assigning values to these variables so that every clause evaluates to 1 (TRUE). Since our goal for this paper is to understand the role of noise in the ability of DMMs to solve the problem, we have generated instances following Ref. [23] so that we can guarantee that they have solutions.
A DMM designed to solve such a problem would use some type of dissipative dynamical systems, where solutions to the 3-SAT correspond to stable equilibrium points of the dynamics (see, e.g., Ref. [1]). However, the choice of the particular dynamical system representing such a problem is not unique. Here, we will consider the one reported in [7, 24] which has been shown to be able to tackle hard instances quite efficiently. These particular DMMs have ODEs of the form (the arguments $v_{n},v_{j},$ and $v_{k}$ are variables appearing in a single clause of the 3-SAT instance):
$\dot{v}_{n}$ $=$ $\sum\limits_{m=1}^{M}x_{l,m}x_{s,m}G_{n,m}(v_{n},v_{j},v_{k})+\left(1+\zeta x_{l,m}\right)\cdot$ (1)
$\left(1-x_{s,m}\right)R_{n,m}(v_{n},v_{j},v_{k}),\qquad\qquad\;n=1,\ldots,N,$
$\dot{x}_{s,m}$ $=$ $\beta\left(x_{s,m}+\epsilon\right)\left(C_{m}(v_{i},v_{j},v_{k})-\gamma\right),\qquad m=1,\ldots,M,$ (2)
$\dot{x}_{l,m}$ $=$ $\alpha\left(C_{m}(v_{i},v_{j},v_{k})-\delta\right),\qquad\qquad\qquad\;\;m=1,\ldots,M,$ (3)
$G_{n,m}$ $=$ $\frac{1}{2}q_{n,m}\text{min}\left[\left(1-q_{j,m}v_{j}\right),\left(1-q_{k,m}v_{k}\right)\right],$ (4)
[…]
$C_{m}$ $=$ $\frac{1}{2}\text{min}\left[\left(1-q_{i,m}v_{i}\right),\left(1-q_{j,m}v_{j}\right),\left(1-q_{k,m}v_{k}\right)\right].$ (6)
Here, $v_{n}\in[-1,1]$ (with $1\leq n\leq N$) are continuous versions of Boolean variables. In an electrical circuit realization of these DMMs, the variables would be, e.g., the voltages at the Boolean gate terminals [1]. Irrespective, we consider $v_{n}>0$ to correspond to TRUE and $v_{n}<0$ to FALSE. The variables, $x_{l,m}$ and $x_{s,m}$ are long and short memory variables bounded in $[1,10^{4}M]$ and $[0,1]$, respectively, whose dynamics are slower than those of the $v_{n}$’s. We set $q_{j,m}=1$ if the literal $j$ enters clause $m$, and $q_{j,m}=-1$ if its negation enters the clause $m$. Additionally, $\alpha$, $\beta$, $\gamma$, $\delta$, $\epsilon$ and $\zeta$ are constants. Following [7], we use the following set of parameter values (tuned for the instances of Ref. [23]): $\alpha=5$, $\beta=20$, $\delta=0.25$, $\gamma=0.05$, $\epsilon=10^{-3}$, $\zeta=0.1$ (for $M/N=7$), and $\zeta=0.001$ (for $M/N=4.3$).
Equations (1)-(3) were integrated numerically using the forward Euler method [25] and added noise to emulate physical noise (see also Secs. 2.33.1 and 3.2 below for more details). The clause function $C_{m}(v_{i},v_{j},v_{k})$ is designed such that $C_{m}<0.5$ if the clause is satisfied. Therefore, the entire instance is solved when $C_{m}<0.5$ for all $m=1,\ldots,M$. It has been shown [7] that for all solvable instances, Eqs. (1)-(3) converge to a solution.
2.2 Lyapunov exponent and power spectrum analyses
Lyapunov exponents – In our study we determine the Lyapunov exponents to characterize the dynamics of DMMs. In the literature, a positive largest Lyapunov exponent (LLE) is typically used as an indicator of chaotic behavior [26]. The usual technique for calculating the LLE is the simultaneous evolution of one perturbation vector along with the system trajectory, with periodic renormalization to keep the norm fixed [27, 28, 29, 30]. However, since the variables in Eqs. (1)-(3) are bounded, this can be problematic because the orientation of maximum expansion may be outside of the phase space. If this occurs, just rescaling the perturbation along this direction would lead to physically invalid states.
To avoid this problem, in the study of numerical noise, we implemented an algorithm introduced by Benettin et al. [31]. The Benettin method avoids this problem by generating a basis of perturbation vectors spanning the tangent space, applying Gram-Schmidt orthonormalization at each step to ensure they remain independent and normalized. This approach eliminates the necessity for reorientation of perturbations, allows the system to select acceptable directions, and maintains all computed exponents physically meaningful in the phase space.
We compute the entire Lyapunov spectrum by slightly perturbing the trajectory in $N$ orthogonal directions with a small perturbation $\Delta\mathbf{v}_{0,k}$ (in the $k$-direction) and integrating each with identical time steps $\Delta t$. We compute the Lyapunov exponents as
$\lambda_{k}=\frac{1}{\Delta t}\text{ln}\left(\frac{||\Delta\mathbf{v}_{k}||}{||\Delta\mathbf{v}_{0,k}||}\right),$ (7)
@@ 2.2 Lyapunov exponent and power spectrum analyses
and choose the largest one (the largest Lyapunov exponent or LLE denoted by $\lambda$) at each time step. At each time step, the perturbations were orthogonalized using the modified Gram-Schmidt orthonormalization process [32, 33] and normalized. In what follows, the following notation is used: $\overline{\lambda}$ for the mean largest Lyapunov exponent (MLLE; time-averaged over one specific trajectory from $t=0$ to the smallest of {solution time, maximum evolution time}), and $\langle\overline{\lambda}\rangle$ for the ensemble averaged mean largest Lyapunov exponent (EAMLLE), where additional averaging occurs over instances of the problem.
@@
In the case of physical noise, the evaluation of the largest Lyapunov exponent involved removing the restriction on the values of $v_{n}$ (refer to the line under Eq. (6)), as discussed in [24] . This removal enabled us to apply the conventional method for calculating the LLE. However, the gradient norm of $v_{n}$ is capped at 10,000 to avoid instability and prevent failure due to unbounded variables and excessively large gradient updates. Additional information can be found in Sec. 3.2 below.
@@
Power spectra – To complement the Lyapunov exponent and have a better understanding of the DMM’s dynamics, we also calculated the power spectrum of the fast variables’ flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t=(\dot{v}_{1},\dot{v}_{2},....\dot{v}_{N})$ to discern the presence of dominant frequencies or a broadband spectrum characteristic of chaos.
@@
For a signal $x(t)$, its power spectrum is usually given by the modulus square of the Fourier components per unit time [34]:
@@
For discrete time signals $x[n]$ of length $N$, the Fast Fourier Transform yields $S[k]=\frac{1}{Nf_{s}}\rvert X[k]\lvert^{2}$, where $X[k]=\sum_{n=0}^{N-1}x[n]e^{-j\pi n/N}$ and $f_{s}=\frac{1}{\Delta t}$ [35]. The power spectrum is obtained by performing this calculation on the right-hand sides of Eq. (1) for each variable independently, and then averaged by the number of variables $N$.
@@
Note that we have opted to consider the power spectrum of the flow vector field, rather than of the $v_{n}(t)$’s to avoid a zero-frequency peak associated with the stable variables’ values ±$\pm 1$. These fast degrees of freedom are the primary dynamic variables whose evolution determines the trajectory in the solution space of the 3-SAT problem. The memory variables, while essential for shaping the dynamics (e.g., inducing transitions or stabilizing solutions), act as “modulators”. Consequently, their time series contain little information about the unstable transitions characteristic of chaos.
2.3 Physical noise

### v2

@@ Abstract
preprint: AIP/123-QED
Developing efficient methods to solve challenging optimization problems is crucial in various technological fields, including logistics, manufacturing, and cryptography. In recent years, significant progress has been made in the area of digital memcomputing – a promising approach to tackle challenging optimization problems. Digital memcomputing machines (DMMs) solve problems by mapping them into dynamical systems and evolving these systems over time. However, in realistic settings, various factors, such as inaccuracies during numerical simulations or imperfections in hardware, introduce sources of noise that may impede the ability of DMMs to find solutions. In this paper, we show that with increasing noise strength, such noise sources can lead the system into a chaotic regime, where the DMM is unable to find the solution. Understanding this noise-induced solvable-unsolvable transition may help to create more stable and noise-tolerant dynamical systems designed to tackle complex problems.
I Introduction
In the field of computing, a new class of non-linear (classical) dynamical systems has been recently introduced, designed to solve complex combinatorial optimization problems Traversa and Di Ventra (2017a). These dynamical systems have been named digital memcomputing machines (DMMs) in view of the fact that they map a finite string of symbols to a finite string of symbols while coupling two types of degrees of freedom: “fast ones” representing the variables of the problem to solve, and “slow ones” representing memory degrees of freedom Di Ventra (2022). Since DMMs are classical systems, their ordinary differential equations (ODEs) can be efficiently simulated on our modern computers. In fact, most of the studies concerning DMMs so far have been performed by simulating their ODEs Traversa and Di Ventra (2017b); Traversa et al. (2018); Traversa and Di Ventra (2018); Sheldon et al. (2019); Bearden, Pei, and Di Ventra (2020); Zhang and Di Ventra (2021); Sharp et al. (2023) [^1] Also see, e.g., case studies in Di Ventra (2022) or those reported by the company MemComputing, Inc.: www.memcpu.com., with a few implementations on small and large FPGA boards Nguyen et al. (2023); Nguyen and Pershin (2024).
Despite much work in this fledgling field has focused on understanding the topological aspects of these dynamical systems Di Ventra, Traversa, and Ovchinnikov (2017); Di Ventra and Ovchinnikov (2019), less attention has been devoted to the role of noise, whether in the form of numerical errors introduced in the discretization of the corresponding ODEs, or the physical noise that would naturally appear in a hardware realization of these machines. Numerical noise can be increased during the time discretization process, by increasing the integration time step, resulting in discrete approximations that alter the path of the system in phase space. Physical noise occurs due to variations in the environment or control parameters, adding complexity and unpredictability. Although there have been multiple studies investigating noise effects in discrete maps and Hamiltonian systems Moss and McClintock (1989); Nychka et al. (1992), the influence of noise on the transition to chaos within continuous-time dynamical systems remains less examined Ovchinnikov and Di Ventra (2019). Nevertheless, the phenomenon of noise-induced chaos has been recognized for several decades. Early works Iansiti et al. (1985); Bulsara et al. (1990) revealed how noise can drive trajectories into fractal basin boundaries, producing switching-like behavior. More recently Lai et al. (2003); Tél et al. (2008), it has been shown that transient chaos, which is present in noise-free systems can, under sufficiently strong noise, become sustained chaos, thus giving rise to chaotic attractors.
Here, we examine the noise-induced transition from a solvable to an unsolvable phase in prototypical DMMs as those introduced in Bearden, Pei, and Di Ventra (2020). Such DMMs have been designed to solve 3-SAT instances (a collection of logical disjunctions of exactly three variables or their negations, related to each other by logical conjunctions), a well-known combinatorial problem Garey and Johnson (1990). These DMMs solve these instances efficiently thanks to the emergence of dynamical long-range order (DLRO) Di Ventra and Ovchinnikov (2019) and the fact that, if properly designed, they admit only saddle points and equilibria corresponding to the solution(s) of the problem instance. No other local minima are present in the phase space.
In the language of field theory, the DLRO of DMMs is expressed by the collective behavior of specific trajectories in the phase space, called instantons, or avalanches Di Ventra and Ovchinnikov (2019), which connect critical points of the flow vector field of the dynamics with decreasing index, namely from saddle points with a certain number of unstable directions to saddle points with less unstable directions. We refer to the book Di Ventra (2022) for a thorough explanation of these statements and to Ref. Bearden, Pei, and Di Ventra (2020) for their mathematical demonstration. The very fact that DMMs employ critical points (saddle points) and trajectories between them (instantons) make them topologically robust, in the sense that their dynamics are fairly insensitive to perturbations. However, we expect this robustness to reach a limit as a function of the noise strength introduced during dynamics, whether numerical or physical.
In the context of nonlinear dynamics, the term “transient chaos” traditionally refers to chaotic behavior that persists only for a finite time before the system evolves towards a fixed point or a limit cycle Lai and Tél (2011). This behavior arises due to the presence of a non-attracting chaotic invariant set, chaotic saddles, in phase space, near which the trajectories evolve chaotically before escaping. However, our focus is on the trajectory that leads to the solution, and we stop the dynamics once the solution is reached and do not analyze the subsequent asymptotic regime. In our system, we can visualize the existence of transient chaos by examining the number of steps until solution as a function of the initial condition (we show this explicitly in the Appendix A). The sensitive non-monotonic dependence we observe is analogous to the results that show the characteristic signature of transient chaos in Ref. Lai and Tél (2011).
The structure of this paper is as follows. Sec. II describes the methods employed in this investigation. In Sec. III, we detail the modeling results, beginning with the case of numerical noise (Sec. III.1) and subsequently addressing the case of physical noise (Sec. III.2). We discuss our results in Sec. IV, and conclude in Sec. V.
II Methods
II.1 DMMs equations for 3-SAT
Let us consider a Boolean satisfiability problem referred to as 3-SAT, which comprises $N$ Boolean variables and $M$ clauses. Each clause is a disjunction of three literals (variables or their negation). Formally, a clause $m$ can be defined by a set $\{\{p^{m}_{1},q^{m}_{1}\},\{p^{m}_{2},q^{m}_{2}\},\{p^{m}_{3},q^{m}_{3}\}\}$, where $p^{m}_{i}$ is the index of the variable entering the clause $m$ at position $i=1,2,3$, and $q^{m}_{i}=\mp 1$ defines whether the variable $\textnormal{v}_{p^{m}_{i}}$ is negated or not, respectively. A solution to the 3-SAT instance involves assigning values to these variables so that every clause evaluates to 1 (TRUE). Since our goal for this paper is to understand the role of noise in the ability of DMMs to solve the problem, we have generated instances following Ref. Barthel et al. (2002) so that we can guarantee that they have solutions.
A DMM designed to solve such a problem would use some type of dissipative dynamical systems, where solutions to the 3-SAT correspond to stable equilibrium points of the dynamics (see, e.g., Ref. Traversa and Di Ventra (2017a)). However, the choice of the particular dynamical system representing such a problem is not unique. Here, we will consider the one reported in Bearden, Pei, and Di Ventra (2020); Pershin and Nguyen (2024) which has been shown to be able to tackle hard instances quite efficiently. These particular DMMs have ODEs of the form (the arguments $\textrm{v}_{n},\textrm{v}_{j},$ and $\textrm{v}_{k}$ are variables appearing in a single clause of the 3-SAT instance):
$\dot{\textnormal{v}}_{n}$ $=$ $\sum\limits_{m=1}^{M}y_{m}x_{m}G_{n,m}+\left(1+\zeta y_{m}\right)\left(1-x_{m}\right)R_{n,m},\hskip 88.2037ptn=1,\ldots,N,$ (1)
$\dot{x}_{m}$ $=$ $\beta\left(x_{m}+\epsilon\right)\left(C_{m}-\gamma\right),\hskip 183.52019ptm=1,\ldots,M,$ (2)
$\dot{y}_{m}$ $=$ $\alpha\left(C_{m}-\delta\right),\hskip 221.93158ptm=1,\ldots,M,$ (3)
$C_{m}$ $=$ $\frac{1}{2}\text{min}\left[\left(1-q^{m}_{1}\textnormal{v}_{p^{m}_{1}}\right),\left(1-q^{m}_{2}\textnormal{v}_{p^{m}_{2}}\right),\left(1-q^{m}_{3}\textnormal{v}_{p_{3}^{m}}\right)\right],$ (4)
$G_{n,m}$ $=$ $\begin{cases}0,\hskip 19.91692ptn\notin\{p^{m}_{1},p^{m}_{2},p^{m}_{3}\},\\ \frac{1}{2}q^{m}_{i}\text{min}\left[\left(1-q^{m}_{j}\textnormal{v}_{p^{m}_{j}}\right),\left(1-q^{m}_{k}\textnormal{v}_{p^{m}_{k}}\right)\right]\mid\{p^{m}_{i}=n,j\neq k\neq i\},\hskip 28.45274pt\text{otherwise},\end{cases}$ (5)
$R_{n,m}$ $=$ $\begin{cases}q^{m}_{i}C_{m}\hskip 28.45274pt\text{if }C_{m}=\frac{1}{2}\left(1-q^{m}_{i}\textnormal{v}_{p^{m}_{i}}\right)\mid\{p^{m}_{i}=n\}\\ 0,\hskip 19.91692pt\text{otherwise}.\end{cases}$ (6)
Here, $N$ is the number of variables, $M$ the number of clauses and $\textnormal{v}_{n}\in[-1,1]$ (with $1\leq n\leq N$) are continuous versions of Boolean variables. In an electrical circuit realization of these DMMs, the variables would be, e.g., the voltages at the Boolean gate terminals Traversa and Di Ventra (2017a). Irrespective, we consider $\textnormal{v}_{n}>0$ to correspond to TRUE and $\textnormal{v}_{n}<0$ to FALSE. The variables, $y_{m}$ and $x_{m}$ are long- and short-memory variables bounded in $1,10^{4}M]$ and $[0,1]$, respectively, whose dynamics are slower than those of the $\textnormal{v}_{n}$’s. The upper bound of $y_{m}$ was arbitrarily imposed to guarantee compactness of the phase space. This bound is high enough to prevent the long-term memory from reaching that ceiling. Additionally, $\alpha$, $\beta$, $\gamma$, $\delta$, $\epsilon$ and $\zeta$ are constants. Following Bearden, Pei, and Di Ventra (2020), we use the following set of parameter values (tuned for the instances of Ref. Barthel et al. (2002)): $\alpha=5$, $\beta=20$, $\delta=0.25$, $\gamma=0.05$, $\epsilon=10^{-3}$, $\zeta=0.1$ (for $M/N=7$), and $\zeta=0.001$ (for $M/N=4.3$).
Equations (1)-(3) were integrated numerically using the forward Euler method Press et al. (2007) and added noise to emulate physical noise (see also Secs. II.3III.1 and III.2 below for more details). The clause function $C_{m}(\textnormal{v}_{p_{1}^{m}},\textnormal{v}_{p_{2}^{m}},\textnormal{v}_{p_{3}^{m}})$ is designed such that $C_{m}<0.5$ if the clause is satisfied. Therefore, the entire instance is solved when $C_{m}<0.5$ for all $m=1,\ldots,M$. It has been shown Bearden, Pei, and Di Ventra (2020) that for all solvable instances, Eqs. (1)-(3) converge to a solution.
II.2 Lyapunov exponent and power spectrum analyses
Lyapunov exponents – In our study we determine the Lyapunov exponents to characterize the dynamics of DMMs. In the literature, a positive largest Lyapunov exponent (LLE) is typically used as an indicator of chaotic behavior Hilborn (2001). The usual technique for calculating the LLE is the simultaneous evolution of one perturbation vector along with the system trajectory, with periodic renormalization to keep the norm fixed Ispolatov et al. (2015); Hochstetter et al. (2021); Aurell et al. (1997); Ercsey-Ravasz and Toroczkai (2011). However, since the variables in Eqs. (1)-(3) are bounded, this can be problematic because the orientation of maximum expansion may be outside of the phase space. If this occurs, just rescaling the perturbation along this direction would lead to unphysical states.
To avoid this problem, in the study of numerical noise, we implemented an algorithm introduced by Benettin et al. Benettin et al. (1980). The Benettin method avoids this problem by generating a basis of perturbation vectors spanning the tangent space, applying Gram-Schmidt orthonormalization at each step to ensure they remain independent and normalized. This approach eliminates the necessity for reorientation of perturbations, allows the system to select acceptable directions, and maintains all computed exponents physically meaningful in the phase space.
We compute the entire Lyapunov spectrum by slightly perturbing the trajectory in $N$ orthogonal directions with a small perturbation $\Delta\mathbf{\textrm{v}}_{0,k}$ (in the $k$-direction) and integrating each with identical time steps $\Delta t$. We compute the instantaneous Lyapunov exponents as
@@ II.2 Lyapunov exponent and power spectrum analyses
where $\|\cdot\|$ denotes the Euclidean norm of the perturbation vector . At each time step, the perturbations were orthogonalized using the modified Gram-Schmidt orthonormalization process Björck and Paige (1992); Björck (1994) and normalized. One then chooses the largest instantaneous Lyapunov exponent denoted by $\lambda$ at each time step. In what follows, the following notation is used: $\overline{\lambda}$ for the mean largest Lyapunov exponent (MLLE; time-averaged over one specific trajectory from $t=0$ to the smallest of {solution time, maximum evolution time}), and $\langle\overline{\lambda}\rangle$ for the ensemble averaged mean largest Lyapunov exponent (EAMLLE), where additional averaging occurs over instances of the problem.
We note that in the noise-free case the system always converges to the point attractors corresponding to the system solutions, and therefore the asymptotic LLE is zero. In this work, we simulated the equations and computed LLE up to the time when the solution was found and thus should be interpreted as finite time LLE. For sufficiently long times, especially in the weak noise regime, LLE would of course decay to zero. This distinction is consistent with the standard definition of noise-induced chaos Lai et al. (2003); Lai and Tél (2011), where the long term exponent remains zero below a critical noise strength and becomes positive only above it.
@@
In the case of physical noise, the evaluation of the largest Lyapunov exponent involved removing the restriction on the values of $\textnormal{v}_{n}$ (refer to the line under Eq. (4)), as discussed in Pershin and Nguyen (2024). This removal enabled us to apply the conventional method for calculating the LLE. However, the gradient norm of $\textnormal{v}_{n}$ is capped at $10^{4}$ to avoid instability and prevent failure due to unbounded variables and excessively large gradient updates. Additional information can be found in Sec. III.2 below.
@@
Power spectra – To complement the Lyapunov exponent and have a better understanding of the DMM’s dynamics, we also calculated the power spectrum of the fast variables’ flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t=(\dot{\textnormal{v}}_{1},\dot{\textnormal{v}}_{2},....\dot{\textnormal{v}}_{N})$ to discern the presence of dominant frequencies or a broadband spectrum characteristic of chaos.
@@
For a signal $x(t)$, its power spectrum is usually given by the modulus square of the Fourier components per unit time van Kampen (1992):
@@
For discrete time signals $x[n]$ of length $N$, the Fast Fourier Transform yields $S[k]=\frac{1}{Nf_{s}}\rvert X[k]\lvert^{2}$, where $X[k]=\sum_{n=0}^{N-1}x[n]e^{-j\pi n/N}$ and $f_{s}=\frac{1}{\Delta t}$ Valsakumar, Satyanarayana, and Sridhar (1997). The power spectrum is obtained by performing this calculation on the right-hand side of Eq. (1) for each variable independently, and then averaged by the number of variables $N$. Therefore $x(t)$ in Eq. (8) would represent here a component of $\textnormal{d}\mathbf{v}/\textnormal{d}t$.
@@
Note that we have opted to consider the power spectrum of the flow vector field, rather than of the $\textnormal{v}_{n}(t)$’s to avoid a zero-frequency peak associated with the stable variables’ values $\pm 1$. These fast degrees of freedom are the primary dynamic variables whose evolution determines the trajectory in the solution space of the 3-SAT problem. The memory variables, while essential for shaping the dynamics (e.g., inducing transitions or stabilizing solutions), act as “modulators”. Consequently, their time series contain little information about the unstable transitions characteristic of chaos.
II.3 Physical noise